Table 1: How to interpret logged models, table adapted from Baileys textbook model equation Log-linear In Y; = Bo + BiX; + eTable 3: Simpsons Variable Descriptions NAME WEIGHT DOUGHNTS FEMALE 0 1 0 0 0 Homer Marge Lisa Bart Comic Book Guy Mr. BurnsTable 6 presents estimates of regression models. Column on contains a regression of WEIGHT on FEMALE. This is a bivariate regTable 6: Regression of Doughnuts on Weight, Linear and Nonlinear Models Dependent variable: WEIGHT log WEIGHT (2) (3) (4) (5)1. regress WEIGHT on FEMALE; this allows us to carry out a difference in means tests that females and males have the same wei1. Interpret the results in Table 6: Regression of Doughnuts on Weight, Linear and Non- linear Models. Use a spreadsheet to p
Answer #1

Q1

a) In colum on we see the results of regression equation where Weight is regressed on Doughnuts. It is a linear regression euqation. In this equation, if we say that weekly the consuption of dounts is increased by ! extra donut, then the weight of the person will increase by 9.224 pounds.

b) In column 2, Weight is regressed on Doughnuts and square of doughnuts. It is a plynomial regression. The weight on consumption of 5 doughnuts is- (5.143*5+0.221*(5^2)= 31.24) 31.24 pounds. In case of consumption of 6 doughnuts (5.143*6+0.221*(6^2)= 38.814) the weight is increased by (38.814-31.24) 7.574 pounts.

The weight on consumption of 19 doughnuts is- (5.143*19+0.221*(19^2)=177.498 ) 177.498 pounds. In case of consumption of 20 doughnuts (5.143*20+0.221*(20^2)= 209.824) the weight is increased by (209.824-177.498) 32.33 pounts.

c) In Table 1, under column 3, Weight is regressed on Log of Doughnuts. It is a linear log regression, thus if the consumption of doughnuts increase by 1% then the weight of the person increases by 0.52 pounds.

d) Column 4 shows the a log linear regression where Log of weight is regressed in Doughnuts. If the consumption of dounts increase by 1 dounut per week then the weight of the person increase by 4.9%.

Q2

a) In Column 2, Weight is the dependent variable and Female and Doughnuts are the independent variables. It shows the regression of Weights on Female dummy and Doughnuts. If the weekly consuption of doughnuts is increased by one-unit then the of the person increases by 8.670 pounds.

b) In Column 3, if the male doughnut consuption is scene, then the dummy variable of FEMALE and interaction variable of FEMALE*DOUGHNUT both will be zero. In such a case, one unit increase in the weekly consumption of dughnut by males will cause their weight to increase by 8.616 pounds.

However if there is a female who is taken into consideration, then if the consumption of doughnut is increased by one-unit per week then the weight is affected in the following manner- (-26.789+8.616+1.276= -16.897). It means that the weight if decreased by 16.9 pounds.